# Coverage and Bias of Street View Imagery in Mapping the Urban Environment

### v2

@@ 5.2.1 Comparing Different SVI Coverage Measurement Methods
Figure 8: Distribution of different SVI coverage completeness metrics and the hot-spot analysis results.
@@ 5.2.2 Built Environment Factors Impacting SVI Coverage
Figure 9: Distribution of built-environment related indicators for local hot-spots and cold-spots of SVI coverage on road length and building facades.
@@ 5.3.1 Robustness of SVI Coverage Across Different SVI Collection Intervals
Figure 10: Box plots drawn based on SVI coverage indicators of different local grids and at different SVI collection intervals.
@@ 5.3.2 Optimal SVI Collection Intervals for Frequency and Completeness
Figure 11: Left: Normalized values of CoC-B and FoC-B of different local grids and at different SVI collection intervals, with the mean values highlighted. Right: A comparison of curve functions fitted with generalized additive model and other modeling methods, based on the mean values.
@@
Figure 12: Derivative curves of completeness and frequency across different local grids and under different isovist analysis radii, and the intersection points between the paired curves.
Figure 13: Spatial and data distribution of optimal SVI collection intervals detected from each H3 level-9 grid in the case study area.
@@ 6.2 Element-level Coverage as A New Dimension for SVI Data Quality
In Experiment 2, the study compares the proposed element-level SVI coverage estimation method, which considers both SVI locations and environmental obstructions, with traditional methods that only consider SVI locations. It was found that, even in regions where traditional methods aim to achieve complete SVI coverage, CoC-A and CoC-B indicators are often low, reflecting a lack of building instances reached by SVI and incomplete facade coverage. The spatial distribution of SVI building facade coverage shows strong spatial auto-correlation, with hot-spots and cold-spots linked to factors like building size, road network centrality, density, and obstacles like greenery and human. The results further explain the implicit prioritization of certain environmental elements in SVI. It is not determined by the subjective intentions of researchers, but rather by the complex spatial configuration and obstruction relations shaped collectively by the roads, buildings and other environmental elements.
@@ 7 Conclusion
This paper introduces a comprehensive workflow to estimate the element-level coverage of SVI, taking urban building facades as an example, and have further explored the potential of coverage extent as novel indicators in validating the usage of SVI in urban and spatial analytics. Our study shows that, despite dense availability on urban road networks, SVI only reaches 62.4% of buildings in the case study area. The completeness of SVI coverage on building facades remains low, averaging 12.4%, with large differences depending on building types and sizes. Further, besides revealing data gaps and inconsistent coverage, our results indicate questionable representativeness of SVI — there are biases in the information collected from SVI, with some instances being over- or under-represented. These biases can impact the integrity of urban studies relying on SVI. For example, if one is using SVI to infer the share of certain building types in streets, with some instances being omitted more often than others in the images collected from cars on roads, the result would not be entirely accurate.
A potential range of optimal SVI sampling intervals, 50-60 m, is identified to help achieve a better application of SVI data. SVI data has been used widely across multiple disciplines, but data quality and integrity have not been given adequate attention. Regarding both the infrastructural and human aspect, for the first time, we reveal at a very high resolution and large-scale, the reach and usability of SVI for urban sensing and mapping. Our study argues that the element-level coverage of SVI, with respect to building, greenery, and other useful street view elements and visual information, should be included as a new dimension for SVI data quality assessment.
@@ Appendix C Summary of the Completeness and Frequency Indicators
The study proposes a indicator system to describe the extent of SVI coverage on environmental features. Below is a table summarizing the SVI coverage indicators proposed and applied in the study.
@@ Appendix D A Re-classification of OSM Building Types
Residential ’apartments’, ’flats’, ’house’, ’terrace’, ’detached’, ’semidetached_house’, ’dormitory’, ’hall_of_residence’, ’cottage’, ’bungalow’, ’terrace_house’, ’council_flats’, ’farm_auxiliary’, ’farm’, ’houseboat’, ’stable’, ’cabin’, ’terraced_house’, ’Nursery,_School’, ’yes;dormitory’
Mixed Use ’yes, office, shop, r’, ’apartments;residenti’, ’apartments;yes’, ’commercial;detached’, ’retail;yes’
Industry and Business ’office’, ’data_center’, ’commercial’, ’warehouse’, ’industrial’, ’light_industrial’, ’factory’, ’manufacture’, ’office;yes’, ’telecommunication’, ’business’, ’artists_studio’
Community Services ’church’, ’university’, ’school’, ’government’, ’public’, ’hospital’, ’college’, ’Community_Building’, ’kindergarten’, ’memorial’, ’student_residence’, ’gatehouse’, ’cafe’, ’greenhouse’, ’monument’, ’pavilion’, ’palace’, ’mosque’, ’synagogue’, ’police_station’, ’religious’, ’clock_tower’, ’village_hall’, ’conservatory’, ’chapel’
Retail ’retail’, ’pub’, ’kiosk’, ’stall’, ’bar’, ’shop’
Transport ’train_station’, ’transportation’, ’ship’, ’boat’, ’bridge’, ’railway_arch’, ’railway’, ’bus’, ’viaduct’, ’tunnel_mouth’, ’tunnel_entrance’, ’bus_garage’
Recreation and Leisure ’civic’, ’hall’, ’ruins’, ’stadium’, ’recreational’, ’gallery’, ’theatre’, ’cinema’, ’museum’, ’sports_centre’, ’sports_hall’, ’swimming_pool’, ’parking’, ’yes;public;sports_ce’
Utilities and Infrastructure ’service’, ’construction’, ’roof’, ’vent_shaft’, ’air_shaft’, ’ventilation_shaft’, ’electricity’, ’substation’, ’gasometer’, ’air_vent’, ’tunnel_shaft’, ’water’
Vacant and Derelict ’vacant’, ’disused_station’, ’abandoned’, ’ruins’
Defence ’guardhouse’, ’bunker’, ’barracks’
Unclassified ’None’, ’yes’, ’no’, ’multiple’, ’part’
@@ Appendix F Completeness Distribution By Building Functions and Sizes
For buildings with different type and size labels, Table 4 compares how the proportions of SVI covered buildings vary. Approximately 62.39% of the total buildings are visible from at least one SVI location within the case study area.
@@ Appendix G Estimation of SVI Covered Population
The research aims to understand to what extent the SVI coverage of residential buildings also achieves coverage of the total population within the study area. The study uses the proportion of residential buildings visible in the SVI at the H3 level-9 grid as a rough representation of the SVI coverage rate of the population in each grid. At the same time, Meta 30 m precision population data is used to estimate the number of people in each H3 grid. The proportion of total population can be covered in SVI can be calculated based on the Formula 7. Based on the proportion of residential buildings covered, it is estimated that approximately 66.2% of the total population in the study area can be covered by SVI.
@@
Acknowledgments
We thank our colleagues at the NUS Urban Analytics Lab for the discussions. The first author is supported by the National University of Singapore under the President Graduate Fellowship. This research is part of the project Large-scale 3D Geospatial Data for Urban Analytics, which is supported by the National University of Singapore under the Start Up Grant R-295-000-171-133.
Author Contributions

### v3

@@ 5.2.1 Comparing different SVI coverage measurement methods
Fig. 8: Distribution of different SVI coverage completeness metrics and the hot-spot analysis results.
@@ 5.2.2 Built environment factors impacting SVI coverage
Fig. 9: Distribution of built-environment related indicators for local hot-spots and cold-spots of SVI coverage on road length and building facades.
@@ 5.3.1 Robustness of SVI coverage across different SVI collection intervals
Fig. 10: Box plots drawn based on SVI coverage indicators of different local grids and at different SVI collection intervals.
@@ 5.3.2 Optimal SVI collection intervals for frequency and completeness
Fig. 11: Left: Normalized values of CoC-B and FoC-B of different local grids and at different SVI collection intervals, with the mean values highlighted. Right: A comparison of curve functions fitted with generalized additive model and other modeling methods, based on the mean values.
@@
Fig. 12: Derivative curves of completeness and frequency across different local grids and under different isovist analysis radii, and the intersection points between the paired curves.
Fig. 13: Spatial and data distribution of optimal SVI collection intervals detected from each H3 level-9 grid in the case study area.
@@ 6.2 Element-level coverage as a new Dimension for SVI data quality
In Experiment 2, the study compares the proposed element-level SVI coverage estimation method, which considers both SVI locations and environmental obstructions, with traditional methods that only consider SVI locations. It is found that, even in regions where traditional methods aim to achieve complete SVI coverage, CoC-A and CoC-B indicators are often low, reflecting a lack of building instances reached by SVI and incomplete facade coverage. The spatial distribution of SVI building facade coverage shows strong spatial auto-correlation, with hot-spots and cold-spots linked to factors like building size, road network centrality, density, and obstacles like greenery and human. The results further explain the implicit prioritization of certain environmental elements in SVI. It is not determined by the subjective intentions of researchers, but rather by the complex spatial configuration and obstruction relations shaped collectively by the roads, buildings and other environmental elements.
@@ 7 Conclusion
This paper introduces a comprehensive workflow to estimate the element-level coverage of SVI, taking urban building facades as an example, and has further explored the potential of coverage extent as novel indicators in validating the usage of SVI in urban and spatial analytics. Our study shows that, despite dense availability on urban road networks, SVI only reaches 62.4 % of buildings in the case study area. The completeness of SVI coverage on building facades remains low, averaging 12.4 %, with large differences depending on building types and sizes. Further, besides revealing data gaps and inconsistent coverage, our results indicate questionable representativeness of SVI — there are biases in the information collected from SVI, with some instances being over- or under-represented. These biases can impact the integrity of urban studies relying on SVI. For example, if one is using SVI to infer the share of certain building types in streets, with some instances being omitted more often than others in the images collected from cars on roads, the result would not be entirely accurate.
A potential range of optimal SVI sampling intervals, 50–60 m, is identified to help achieve a better application of SVI data. SVI data has been used widely across multiple disciplines, but data quality and integrity have not been given adequate attention. Regarding both the infrastructural and human aspect, for the first time, we reveal at a very high resolution and large-scale, the reach and usability of SVI for urban sensing and mapping. Our study argues that the element-level coverage of SVI, with respect to building, greenery, and other useful street view elements and visual information, should be included as a new dimension for SVI data quality assessment.
@@
Acknowledgments
We thank our colleagues at the NUS Urban Analytics Lab for the discussions. The first author is supported by the National University of Singapore under the President’s Graduate Fellowship. This research is part of the project Large-scale 3D Geospatial Data for Urban Analytics, which is supported by the National University of Singapore under the Start Up Grant R-295-000-171-133.
Author contributions
@@ Appendix C Summary of the Completeness and Frequency Indicators
The study proposes an indicator system to describe the extent of SVI coverage on building element. Below is a table summarizing the SVI coverage indicators proposed and applied in the study.
@@ Appendix D A Re-classification of OSM Building Types
Residential ‘apartments’, ‘flats’, ‘house’, ‘terrace’, ‘detached’, ‘semidetached_house’, ‘dormitory’, ‘hall_of_residence’, ‘cottage’, ‘bungalow’, ‘terrace_house’, ‘council_flats’, ‘farm_auxiliary’, ‘farm’, ‘houseboat’, ‘stable’, ‘cabin’, ‘terraced_house’, ‘Nursery,_School’, ‘yes;dormitory’
Mixed Use ‘yes, office, shop, r’, ‘apartments;residenti’, ‘apartments;yes’, ‘commercial;detached’, ‘retail;yes’
Industry and Business ‘office’, ‘data_center’, ‘commercial’, ‘warehouse’, ‘industrial’, ‘light_industrial’, ‘factory’, ‘manufacture’, ‘office;yes’, ‘telecommunication’, ‘business’, ‘artists_studio’
Community Services ‘church’, ‘university’, ‘school’, ‘government’, ‘public’, ‘hospital’, ‘college’, ‘Community_Building’, ‘kindergarten’, ‘memorial’, ‘student_residence’, ‘gatehouse’, ‘cafe’, ‘greenhouse’, ‘monument’, ‘pavilion’, ‘palace’, ‘mosque’, ‘synagogue’, ‘police_station’, ‘religious’, ‘clock_tower’, ‘village_hall’, ‘conservatory’, ‘chapel’
Retail ‘retail’, ‘pub’, ‘kiosk’, ‘stall’, ‘bar’, ‘shop’
Transport ‘train_station’, ‘transportation’, ‘ship’, ‘boat’, ‘bridge’, ‘railway_arch’, ‘railway’, ‘bus’, ‘viaduct’, ‘tunnel_mouth’, ‘tunnel_entrance’, ‘bus_garage’
Recreation and Leisure ‘civic’, ‘hall’, ‘ruins’, ‘stadium’, ‘recreational’, ‘gallery’, ‘theatre’, ‘cinema’, ‘museum’, ‘sports_centre’, ‘sports_hall’, ‘swimming_pool’, ‘parking’, ‘yes;public;sports_ce’
Utilities and Infrastructure ‘service’, ‘construction’, ‘roof’, ‘vent_shaft’, ‘air_shaft’, ‘ventilation_shaft’, ‘electricity’, ‘substation’, ‘gasometer’, ‘air_vent’, ‘tunnel_shaft’, ‘water’
Vacant and Derelict ‘vacant’, ‘disused_station’, ‘abandoned’, ‘ruins’
Defence ‘guardhouse’, ‘bunker’, ‘barracks’
Unclassified ‘None’, ‘yes’, ‘no’, ‘multiple’, ‘part’
@@ Appendix F Completeness Distribution By Building Functions and Sizes
For buildings with different type and size labels, Table 4 compares how the proportions of SVI covered buildings vary. Approximately 62.4 % of the total buildings are visible from at least one SVI location within the case study area.
@@ Appendix G Estimation of SVI Covered Population
The research aims to understand to what extent the SVI coverage of residential buildings also achieves coverage of the total population within the study area. The study uses the proportion of residential buildings visible in the SVI at the H3 level-9 grid as a rough representation of the SVI coverage rate of the population in each grid. At the same time, Meta 30 m precision population data is used to estimate the number of people in each H3 grid. The proportion of total population covered in SVI can be calculated based on the Formula 7. Based on the proportion of residential buildings covered, it is estimated that approximately 66.2 % of the total population in the study area can be covered by SVI.